# Safe Learning for Contact-Rich Robot Tasks: A Survey from Classical Learning-Based Methods to Safe Foundation Models

### v1

@@ 3.6.1 Safety, Efficiency and Task Objectives
Safety is usually assessed through the violation of predefined constraints [13, 11] or the value of a cost function [65]. In physical interaction tasks, it is often quantified through physical metrics such as force, acceleration and velocity [36, 44]. Control stability and human safety are important subcategories of the safety objective. Stability can be measured by the convergence rate [21, 2] or the energy flow of the system [zanella2024learning, zhang2025passivity, 375]. The human safety includes primarily the physical safety of the human in direct contact tasks like massaging or bathing assistance [19, 5], however, perceived safety concepts such as trust, predictability and comfort gain importance for the acceptance of robots in social spaces [52]. A relevant work [9] evaluates the perceived safety through 7-point Likert survey in their dressing assistance experiments.
@@ 3.6.2 Trade-off Between Objectives
The three objectives (Fig. 5) may conflict in many instances, thus the designers often analyze the trade-off between these. The trade-off can be measured by a combined metric, such as the ratio of the safe and successful experiments to constraint violating ones [91, zhang2024srl]. A safety-efficiency reward function was proposed by [28] to achieve a good trade-off. On the other hand, [11] and [13] compare the numbers of successful and violating runs side-by-side instead of directly calculating a ratio between them.
Machine learning metrics such as prediction accuracy has been used to measure task objectives, however, it is also relevant in works that use learned models for safety-related predictions [21, 4]. Just as the learning curves provide insights on the success and efficiency trade-off, some works present the learning curves of success, violation and their ratio to portray the trade-off between the three objectives during training [91, zhang2024srl].
Although shifting the focus to safety can decrease the task performance or learning efficiency, these objectives can be aligned in the case of physically interpretable metrics such as force and motion tracking errors, and energy consumption. Physically meaningful results such as the force-velocity behavior can provide intuition towards the safety of the system [1, 56]. Tracking the target force without large deviations in the Lyapunov-based boundaries directly translates to both the control stability and task achievement quality [2]. On the other hand, learning energy-based stability from data directly translates to energy efficiency in deployment [zhang2025passivity].
@@ 3.6.3 Improved Evaluation
The safety aspect of a learning system can be further evaluated through added noise and disturbance. [5] evaluates the convergence of their trajectory tracking controller under external disturbance. The noise can be injected in observation-space [10, 18, 94] or action-space [20]. These tests help assess the limitations and sensing requirements of the systems. [87] evaluate their system under uniformly sampled simulation parameters, in addition to localization uncertainties. This allows demonstrating the robustness and generalization capacity of the method.
@@
Most of the metrics can be presented at different levels of detail. In Fig. 6, we define 3 levels of analyses to establish a common understanding and encourage more insightful results: 1. global(high-level) analyses present aggregated summary statistics such as the maximum force, total number of violations, or mean ($\pm$ variance) tracking error over many episodes. 2. variational(mid-level) analyses show the trends of metrics among many episodes or seeds, such as a learning curve, a histogram of execution times, or a scatter plot of success ratio versus force violation. 3. exemplar(low-level) analyses present descriptive case studies, often real-time measurements from a single episode, such as motion trajectories, force measurements, or even multi-media material. Going downwards, the amount of detail and depth of analysis increase; going upwards, the results get more general and memorable. Variational comparisons allow a more in-depth analysis of the system behavior from temporal, spatial or multi-variate perspective. Exemplar results can give valuable insights on why and how the method works, or when it fails, without deriving general conclusions. We recommend providing analyses at all levels for a more comprehensive communication of the results.
@@ 3.7.2 Low-Level Control
Low-level safety control focuses on ensuring safe interactions at the actuation and feedback level. This includes real-time enforcement of physical constraints such as force, torque, impedance, and stability, which are especially critical in contact-rich scenarios. Learning-based methods here typically involve adaptive impedance control, learning CBFs, model-predictive safety filters, or robust policy adaptation techniques. [30]
Compliant control and impedance learning. A widely adopted strategy is to learn variable impedance or admittance parameters rather than direct torques [23, 25, 28, 317]. By modulating stiffness and damping online based on contact feedback, these methods ensure compliant behavior that naturally limits interaction forces and absorbs impacts [1, 40, 74]. Recent approaches also integrate passivity constraints or energy tanks into the learning process to guarantee coupled stability during dynamic contact [zhang2025passivity, 35].
Safety filters and barriers. To enforce hard safety constraints (e.g., joint limits, friction cones, maximum force), Control Barrier Functions (CBFs) and safety filters are often employed as a protective layer wrapping the learned policy [29, 152, 235]. These mechanisms monitor the control input and minimally deviate it to maintain the system within a forward-invariant safe set. Model Predictive Safety Filters (MPSF) extend this by reasoning over a short horizon to prevent inevitable future violations, which is crucial for systems with inertia or actuation limits [21, 212].
Robust and adaptive safety. For scenarios with significant model uncertainty (e.g., unknown friction or payload), robust control techniques such as Sliding Mode Control (SMC) or adaptive control are combined with learning to maintain safety margins [5, 60, 94]. These low-level controllers compensate for disturbances and unmodeled dynamics that might otherwise destabilize a high-level learned policy.
@@ 3.7.3 End-to-End Safety Enhancement
End-to-end methods aim to integrate perception, decision-making, and actuation within a unified model while embedding safety either implicitly (e.g., through regularization) or explicitly (e.g., via constrained optimization or safety critics). This includes safe reinforcement learning frameworks that learn both task policies and safety constraints simultaneously from raw sensory data [287, 99, 236]. For instance, [100] proposed a neural network-based approach that learns to predict safe actions directly from raw sensory inputs, such as images and proprioceptive data, allowing the robot to adapt its behavior in real-time to avoid unsafe actions during contact-rich tasks.
Recently, Vision-Language-Action (VLA) models have emerged as a powerful paradigm for end-to-end safety enhancement. These models bridge perception, reasoning, and control by learning from large-scale multimodal datasets that include language instructions, visual observations, and robot actions [330, 33, 10705419]. Safety is enhanced in VLA systems through several mechanisms: (i) Safety-aware prompting and grounding: VLA models can be conditioned on safety-specific prompts or rules (e.g., ”move slowly near the glass”) to generate compliant behaviors [8, 51]. (ii) Safety critics and filtering: Auxiliary safety modules or critics can be trained to evaluate the risk of generated actions before execution, filtering out potentially hazardous commands [376, 360, 115]. (iii) Multimodal alignment: By aligning visual and linguistic representations with physical safety constraints (e.g., force limits), VLA models can better generalize safety concepts to novel objects and environments [309, 307, 218]. While these methods promise scalability and generalization, especially in contact-uncertain environments, they often suffer from limited interpretability and pose challenges in formal safety verification compared to modular architectures.
@@ 3.7.4 Hybrid Methods
Plan–learn–filter pipelines.
A high‑level planner (graph, sampling, optimization, or semantic decomposition) proposes nominal contact strategies or intermediate sub‑goals; a learned policy (often variable impedance/admittance or residual) generates parameterized actions (pose, force, stiffness); then a runtime safety layer (CBF / reachability / MPC shield/projection) minimally modifies unsafe commands while preserving progress [25, 28, 29, 26, martinez2015safe, yuan2022safe]. This layering reduces unsafe transients at approach, impact, and sustained contact phases and limits over‑conservatism by deferring exact constraint satisfaction to an online filter.
Compliant action abstraction + certified projection.
Rather than outputting raw torques, hybrid methods learn action abstractions (desired Cartesian displacements, force targets, diagonal or structured stiffness gains) that preserve passivity margins and expose explicit hooks for safety modification [zhang2025passivity, zhang2024srl, 25, 28, 317, roveda2020model]. Certified projection (QP with CBF constraints, shield veto, reachability pre‑checks) enforces forward invariance or constraint satisfaction without retraining the core policy [29, 26, yuan2022safe, 152].
Generative/residual policies under safety wrappers.
Diffusion or residual RL policies produce candidate motions or compliance updates; a safety critic or barrier/MPC layer evaluates predicted force envelopes, energy flow, or constraint margins before execution (“generate–score–project”) [100, zhang2024srl, 29, zhang2025passivity, 362, 117, 12]. Residual admittance and stiffness adaptation compensate for unmodeled contact geometry while the wrapper prevents force spikes [25, 28, 317, 178].
Semantic / VLM–VLA hybrid safety.
Vision–language(-action) and emerging foundation model components parse unstructured instructions and scene context into structured constraint objects (force caps, forbidden surface masks, approach cones) [33, 330, 10705419]. These are handed to the plan–learn–filter stack: semantic planner produces goal / sequence; learned variable impedance or residual policy outputs compliant parameters [376, 360]; runtime CBF / reachability / MPC shield enforces limits; logging layer records violations for calibration [25, 28, 29, yuan2022safe]. Multimodal fusion (vision + force) disambiguates occlusions and anticipates risky contact transitions, while structured prompting and action schema reduce hallucination‑induced unsafe commands [100, 33, 330]. This hybridization aligns high‑level semantic intent with certifiable low‑level execution, enabling scalable yet physically grounded safe manipulation in contact‑rich settings.
@@ 3.8 Safety Enforcement Spaces
In safe learning for robotic contact-rich tasks, a central consideration is where safety constraints are enforced within the control or learning pipeline. Safety can be realized at different levels of the robot’s architecture, from the physical interaction interface (task space), to the internal actuation system (joint space), and even at the policy decision layer. Each of these domains provides distinct opportunities and challenges for incorporating safety. Task-space enforcement directly regulates end-effector motion and contact forces with the environment, joint-space methods shape internal robot behavior to avoid unsafe configurations or torques, while policy-level mechanisms ensure that the agent’s decision-making respects safety constraints from the outset. This section categorizes and compares prior works based on the level at which safety is defined and enforced, highlighting dominant trends, common control strategies, and emerging hybrid approaches.
@@ 3.8.1 Task Space
Many approaches ensure safety directly in the task (Cartesian) space, meaning they constrain the robot’s end-effector motion and interaction forces. These methods typically impose limits on end-effector position, orientation, velocity, or contact force to keep them within safe bounds during contact-rich tasks. For example, some works explicitly bound the forces and torques at the end-effector to prevent excessive contact pressures, integrating such constraints into the controller or learning process [1, 3].
By formulating safety at the task level, the robot only executes motions that respect predefined safe regions and force thresholds in the workspace. A common strategy is to employ impedance or admittance control in Cartesian space, adjusting the end-effector’s stiffness or damping in real time based on feedback to maintain compliant, gentle contact [4, 6, 7, 12, 16]. This allows the robot to absorb shocks and limit forces during manipulation while still following the desired trajectory. Several researchers integrate these variable impedance frameworks with learning algorithms, so the policy itself can modulate compliance for safety, for instance, by outputting stiffness parameters or virtual force targets that keep interactions soft and stable [23, 25, 107].
In reinforcement learning settings, safety constraints in task space are often enforced by penalizing or stopping unsafe behavior. Episodes might terminate or incur a large cost if the robot’s force sensor detects a contact force exceeding a set threshold [10, 20, 62, 84]. Likewise, the action space can be shaped to avoid dangerous moves. Some methods confine the end-effector’s motions to a safe subset of the workspace or around demonstrated trajectories, preventing the agent from exploring configurations that could lead to collisions or high forces [11, 22, 96, 98]. Through these measures, task-space safety approaches directly guard the physical interaction interface of the robot, ensuring that all contact with the environment or humans remains within safe limits.
Beyond basic compliance and force limiting, formal safety techniques have also been applied at the task-space level. Control barrier functions and similar safety filters can constrain the end-effector’s position and force within a known safe region at all times, overriding any controller command that would violate safety limits [14, 57, 80].
Some works derive safe sets from data or models, for instance, learning a set of permissible end-effector states from demonstrations or via Gaussian processes, and then formulating the task-space controller or planner to keep the robot’s trajectory inside this safe latent manifold [21, 102].
There are also planning-based approaches that build safety into motion generation. Before execution or learning, they compute contact-rich trajectories that satisfy contact stability, task unrelated collision avoidance, and joint limits, so that any subsequent policy learning begins with inherently safe trajectories [8, 34].
Energy-based methods add another layer of safety by ensuring the robot behaves as a passive system in task space. For example, adaptive impedance controllers with energy tanks or passivity constraints guarantee that the robot cannot inject unbounded energy into the environment, thus preventing spikes in force or unstable interactions [35, 75].
@@
Notably, in human–robot interaction scenarios, task-space safety often involves maintaining a safe distance and force relative to humans. Techniques like dynamic safety fields enforce minimum separation distances and speed limits for the end-effector when a person is nearby, and controllers can be designed to comply with industry safety standards, e.g. ISO force limits, during physical contact [15, 51, 93].
In summary, by enforcing safety in the task space, these approaches directly regulate the robot’s end-effector behavior through keeping contact forces, motions, and interactions within safe, tolerable bounds to guarantee safe manipulation.
@@ 3.8.2 Joint Space
Another line of research enforces safety in the joint (configuration) space, focusing on the robot’s internal angles, velocities, and torques. Joint-space safety methods ensure the robot’s actuators never produce motions that could be unsafe, thereby indirectly protecting the end-effector and its environment.
One way this is achieved is by augmenting the robot’s low-level controllers with safety constraints. For instance, joint-level control barrier functions can impose limits on joint motion or torque. A controller will automatically adjust or saturate commands so that resulting contact forces do not exceed a strict threshold [44]. Similarly, variable impedance control can be formulated at the joint level with guarantees of stability. By enforcing that the learned joint stiffness and damping matrices remain positive-definite (and hence passive), the system maintains Lyapunov stability and safe behavior throughout contact interactions [53, 59].
Some approaches use adaptive joint-torque controllers that account for external forces. They regulate joint torques in response to contact feedback, ensuring that sudden impacts are absorbed by the compliance in the joints rather than causing dangerous forces [60, 94]. In reinforcement learning contexts where policies output joint commands, researchers incorporate joint-space safety by constraining the action range and optimizing within safe bounds. For example, an RL policy might be allowed to command only small joint movements directly, while larger repositioning actions invoke a motion planner that finds collision-free joint trajectories, thereby avoiding unsafe jumps in configuration [99].
It is also common to include penalties for approaching joint limits or producing high torques in the reward function, so the agent naturally prefers to stay within a safe range of motion [97]. In one case, a policy was trained to perform robotic massage by learning joint-level impedance parameters. To keep the interaction safe with the human, the algorithm limited joint velocities and forces via compliance control, effectively bounding the output torque delivered to the person [46]. By prioritizing safety in the joint space, these methods keep the robot’s fundamental degrees of freedom under control through preventing excessive joint speeds or forces, which in turn helps maintain stable and safe interactions at the task space as well.

### v2

@@ 3.6.1 Safety, Efficiency and Task Objectives
Safety is usually assessed through the violation of predefined constraints [151, 234] or the value of a cost function [366, 90]. In physical interaction tasks, it is often quantified through physical metrics such as force, acceleration and velocity [127, 323]. Control stability and human safety are important subcategories of the safety objective. Stability can be measured by the convergence rate [217, 167] or the energy flow of the system [364, 373, 112]. The human safety includes primarily the physical safety of the human in direct contact tasks like massaging or bathing assistance [339, 345], however, perceived safety concepts such as trust, predictability and comfort gain importance for the acceptance of robots in social spaces [73]. A relevant work [290] evaluates the perceived safety through 7-point Likert survey in their dressing assistance experiments.
@@ 3.6.2 Trade-off Between Objectives
The three objectives (Fig. 7) may conflict in many instances, thus the designers often analyze the trade-off between these. Having both task and safety objectives increases the optimization complexity, and consequently the amount of data to learn effective policies [90]. The trade-off can be measured by a combined metric, such as the ratio of the safe and successful experiments to constraint violating ones [301, 374]. A safety-efficiency reward function was proposed by [169] to achieve a good trade-off. On the other hand, [234], [1] and [151] compare the numbers of successful and violating runs side-by-side instead of directly calculating a ratio between them.
Machine learning metrics such as prediction accuracy has been used to measure task objectives, however, it is also relevant in works that use learned models for safety-related predictions [217, 232]. Just as the learning curves provide insights on the success and efficiency trade-off, some works present the learning curves of success, violation and their ratio to portray the trade-off between the three objectives during training [301, 374].
Although shifting the focus to safety can decrease the task performance or learning efficiency, these objectives can be aligned in the case of physically interpretable metrics such as force and motion tracking errors, and energy consumption. Physically meaningful results such as the force-velocity behavior can provide intuition towards the safety of the system [56, 359]. Tracking the target force without large deviations in the Lyapunov-based boundaries directly translates to both the control stability and task achievement quality [167]. On the other hand, learning energy-based stability from data directly translates to energy efficiency in deployment [373].
@@ 3.6.3 Improved Evaluation
The safety aspect of a learning system can be further evaluated through added noise and disturbance. [345] evaluates the convergence of their trajectory tracking controller under external disturbance. The noise can be injected in observation-space [226, 181, 331] or action-space [242]. Vision-based systems can be tested with changing backgrounds of cluttered scenes [389]. These tests help assess the limitations and sensing requirements of the systems. [281] evaluate their system under uniformly sampled simulation parameters, in addition to localization uncertainties. [45] compares the safety and success of different approaches in non-stationary environments. These and similar evaluation efforts allow demonstrating the robustness and generalization capacity of a method.
@@
Apart from the added conditions, the evaluation can be strengthened by presenting metrics at different levels of detail. In Fig. 8, we define 3 levels of analyses to establish a common understanding and encourage more insightful results: 1. global(high-level) analyses present aggregated summary statistics such as the maximum force, total number of violations, or mean ($\pm$ variance) tracking error over many episodes. 2. variational(mid-level) analyses show the trends of metrics among many episodes or seeds, such as a learning curve, a histogram of execution times, or a scatter plot of success ratio versus force violation. 3. exemplar(low-level) analyses present descriptive case studies, often real-time measurements from a single episode, such as motion trajectories, force measurements, or even multi-media material. Going downwards, the amount of detail and depth of analysis increase; going upwards, the results get more general and memorable. Variational comparisons allow a more in-depth analysis of the system behavior from temporal, spatial or multi-variate perspective. Exemplar results can give valuable insights on why and how the method works, or when it fails, without deriving general conclusions. We recommend providing analyses at all levels for a more comprehensive communication of the results.
@@ 3.7.2 Low-Level Control
Low-level safety control focuses on ensuring safe interactions at the actuation and feedback level. This includes real-time enforcement of physical constraints such as force, torque, impedance, and stability, which are especially critical in contact-rich scenarios. Learning-based methods here typically involve adaptive impedance control, learning CBFs, model-predictive safety filters, or robust policy adaptation techniques [108].
Compliant control and impedance learning A widely adopted strategy is to learn variable impedance or admittance parameters rather than direct torques [350, 351, 169, 222]. By modulating stiffness and damping online based on contact feedback, these methods ensure compliant behavior that naturally limits interaction forces and absorbs impacts [56, 11, 175]. Recent approaches also integrate passivity constraints or energy tanks into the learning process to guarantee coupled stability during dynamic contact [373, 60].
Safety filters and barriers To enforce hard safety constraints (e.g., joint limits, friction cones, maximum force), CBFs and safety filters are often employed as a protective layer wrapping the learned policy [132, 171]. These mechanisms monitor the control input and minimally deviate it to maintain the system within a forward-invariant safe set. Model Predictive Safety Filters (MPSF) extend this by reasoning over a short horizon to prevent inevitable future violations, which is crucial for systems with inertia or actuation limits [217, 15].
Robust and adaptive safety For scenarios with significant model uncertainty (e.g., unknown friction or payload), robust control techniques such as SMC or adaptive control are combined with learning to maintain safety margins [345, 354, 331]. These low-level controllers compensate for disturbances and unmodeled dynamics that might otherwise destabilize a high-level learned policy.
@@ 3.7.3 End-to-End Safety Enhancement
End-to-end methods aim to integrate perception, decision-making, and actuation within a unified model while embedding safety either implicitly (e.g., through regularization) or explicitly (e.g., via constrained optimization or safety critics). This includes safe reinforcement learning frameworks that learn both task policies and safety constraints simultaneously from raw sensory data [48, 347, 277]. For instance, Bahl et al. [14] proposed a neural network-based approach that learns to predict safe actions directly from raw sensory inputs, such as images and proprioceptive data, allowing the robot to adapt its behavior in real-time to avoid unsafe actions during contact-rich tasks.
Recently, VLA models have emerged as a powerful paradigm for end-to-end safety enhancement. These models bridge perception, reasoning, and control by learning from large-scale multimodal datasets that include language instructions, visual observations, and robot actions [307, 367, 388]. Safety is enhanced in VLA systems through several mechanisms: (i) Safety-aware prompting and grounding: VLA models can be conditioned on safety-specific prompts or rules (e.g., “move slowly near the glass”) to generate compliant behaviors [225, 303]. (ii) Safety critics and filtering: Auxiliary safety modules or critics can be trained to evaluate the risk of generated actions before execution, filtering out potentially hazardous commands [370, 21, 355]. (iii) Multimodal alignment: By aligning visual and linguistic representations with physical safety constraints (e.g., force limits), VLA models can better generalize safety concepts to novel objects and environments [329, 200, 75]. While these methods promise scalability and generalization, especially in contact-uncertain environments, they often suffer from limited interpretability and pose challenges in formal safety verification compared to modular architectures.
@@ 3.7.4 Hybrid Methods
Plan-learn-filter pipelines.
A high‑level planner (graph, sampling, optimization, or semantic decomposition) proposes nominal contact strategies or intermediate sub-goals; a learned policy (often variable impedance/admittance or residual) generates parameterized actions (pose, force, stiffness); then a runtime safety layer (CBF / reachability / MPC shield or projection) minimally modifies unsafe commands while preserving progress [351, 169, 132, 348, 206, 362]. This layering reduces unsafe transients at approach, impact, and sustained contact phases by deferring exact constraint satisfaction to an online filter.
Compliant action abstraction and certified projection.
Rather than outputting raw torques, hybrid methods learn action abstractions (desired Cartesian displacements, force targets, diagonal or structured stiffness gains) that preserve passivity margins and expose explicit hooks for safety modification [373, 374, 351, 169, 222, 248]. Certified projection (QP with CBF constraints, shield veto, reachability pre‑checks) enforces forward invariance or constraint satisfaction without retraining the core policy [132, 348, 362, 123].
Generative or residual policies under safety wrappers.
Diffusion or residual RL policies produce candidate motions or compliance updates; a safety critic or barrier/MPC layer evaluates predicted force envelopes, energy flow, or constraint margins before execution (“generate-score-project”) [14, 374, 132, 373, 81, 336, 3]. Residual admittance and stiffness adaptation compensate for unmodeled contact geometry while the wrapper prevents force spikes [351, 169, 222, 244].
Semantic or VLM/VLA hybrid safety.
Vision-language(-action) and emerging foundation model components parse unstructured instructions and scene context into structured constraint objects [367, 307, 388]. These are handed to the plan-learn-filter stack: semantic planner produces goal or sequence; learned variable impedance or residual policy outputs compliant parameters [370, 21]; Runtime CBF, reachability, and MPC shield enforces limits; logging layer records violations for calibration [351, 169, 132, 362]. Multimodal fusion, especially vision and tactile and force, disambiguates occlusions and anticipates risky contact transitions, while structured prompting and action schema reduce hallucination-induced unsafe commands [14, 367, 307]. This hybridization aligns high‑level semantic intent with certifiable low‑level execution, enabling scalable yet physically grounded safe manipulation in contact‑rich settings.
@@ 3.8 Safety Enforcement Spaces
In the context of learning for contact-rich manipulation, safety is not only a monolithic concept but also is enforced hierarchically across the robot’s control architecture. Constraints can be imposed at the level of physical interaction (task space), internal actuation (joint space), or within the decision-making abstraction (policy space). Each of these domains provides distinct opportunities and challenges for incorporating safety. Task-space enforcement directly regulates end-effector motion and contact forces with the environment, joint-space methods shape internal robot behavior to avoid unsafe configurations or torques, while policy-level mechanisms ensure that the agent’s decision-making respects safety constraints from the outset. This section categorizes existing literature based on these architectural levels, synthesizing the dominant control strategies and emerging hybrid paradigms.
@@ 3.8.1 Task Space
Task-space mechanisms operate directly at the interface of physical interaction, regulating the kinematics and dynamics of the end-effector. By defining safety in Cartesian coordinates, these methods directly constrain the contact forces and motions exerted on the environment. A predominant strategy for task-space safety is the use of compliance and stiffness modulation. Rather than tracking rigid trajectories, impedance and admittance control schemes allow the robot to deviate from reference paths in response to external forces, thereby absorbing impact energy and maintaining stable contact [232, 106, 1, 3, 165]. The works in [56, 365] explicitly bound the forces and torques at the end-effector to prevent excessive contact pressures, integrating such constraints into the controller and learning process. Recent learning-based frameworks have extended this architecture by parameterizing variable impedance policies, where the agent learns to output context-aware stiffness or virtual force targets to ensure gentle interaction [350, 351, 205]. To guarantee stability in these interactions, energy-based approaches often enforce passivity constraints, ensuring that the robot behaves as a passive system and does not inject unbounded energy into the environment [60, 131].
Beyond inherent compliance, formal constraints are frequently applied to bound the workspace. CBFs and safety filters are employed to strictly confine end-effector states within a safe set, overriding policy actions that threaten to violate positional or force limits [313, 369, 316]. These safe manifolds can be learned from demonstrations or modeled via Gaussian processes, restricting the planner to regions with high probability of safety [217, 276]. In reinforcement learning settings, task-space safety is often formulated via reward shaping to penalize excessive contact forces [226, 242, 258, 324], or by restricting the action space to a safe subset of the workspace to prevent collisions during exploration[234, 138, 24, 385]. The works in [225, 397] also use planning-based approaches to build safety into motion generation. Before execution or learning, they compute contact-rich trajectories that satisfy contact stability, task unrelated collision avoidance, and joint limits, so that any subsequent policy learning begins with inherently safe trajectories.
@@
Notably, in human–robot interaction scenarios, the constraints are adapted to dynamic environments, utilizing safety fields to enforce separation distances, and controllers are designed to comply with industry safety standards, e.g. ISO force limits, during physical contact [396, 303, 71].
@@ 3.8.2 Joint Space
While task-space methods focus on extrinsic interaction, joint-space enforcement prioritizes the robot’s intrinsic integrity. These methods operate in the configuration space, constraining joint angles, velocities, and torques to prevent self-collisions, singularity proximity, and actuator saturation.
Safety in this domain is often realized by augmenting low-level controllers with limiting functions. Joint-level CBFs and saturation filters ensure that commanded torques do not exceed hardware capabilities or result in unsafe kinetic energy, thereby indirectly limiting the force transfer to the environment [323]. Stability is further ensured through Lyapunov-based formulations, such as enforcing positive-definiteness in learned joint stiffness matrices to guarantee system passivity [64, 236].
To mitigate the risks of dynamic impacts, adaptive joint-torque controllers utilize proprioceptive feedback to dissipate shock through joint compliance [354, 331]. In the context of RL, policies operating in joint space are frequently constrained to prevent erratic high-frequency actuation. Strategies include penalizing proximity to joint limits [32] or employing hierarchical schemes where the RL agent dictates high-level goals while a lower-level motion planner generates smooth, collision-free configuration trajectories [347]. Specific applications, such as robotic massage, have demonstrated that limiting joint-space velocities and applying compliance controls effectively bounds the output torque, ensuring safe physical interaction through intrinsic actuation limits [170].